### v4

@@ Abstract
Deepfake is a technology dedicated to creating highly realistic facial images and videos under specific conditions, which has significant application potential in fields such as entertainment, movie production, and digital human creation, to name a few. With the advancements in deep learning, techniques primarily represented by Variational Autoencoders and Generative Adversarial Networks have achieved impressive generation results. More recently, the emergence of diffusion models with powerful generation capabilities has sparked a renewed wave of research. In addition to deepfake generation, corresponding detection technologies continuously evolve to regulate the potential misuse of deepfakes, such as privacy invasion and phishing attacks. This survey comprehensively reviews the latest developments in deepfake generation and detection, summarizing and analyzing current state-of-the-arts in this rapidly evolving field. First, we unify task definitions, comprehensively introduce datasets and metrics, and discuss developing technologies. Then, we discuss the development of several related sub-fields and focus on researching four representative deepfake fields: face swapping, face reenactment, talking face generation, and facial attribute editing, as well as forgery detection. Subsequently, we comprehensively benchmark representative methods on popular datasets for each field, fully evaluating the latest and most influential published works. Finally, we analyze the challenges and future research directions of the discussed fields. We closely follow the latest developments in this project.
Keywords:
@@
Deepfake Generation Face Swapping Face Reenactment Talking Face Generation Facial Attribute Editing Forgery detection Survey
1 Introduction
@@
Artificial Intelligence Generated Content (AIGC) garners considerable attention sha2023deep in academia and industry. Deepfake generation, as one of the important technologies in the generative domain, gains significant attention due to its ability to create highly realistic facial media content. This technique transitions from traditional graphics-based methods to deep learning-based approaches. Early methods employ advanced Variational Autoencoder kingma2014auto ; sohn2015learning ; van2017neural (VAE) and Generative Adversarial Network (GAN) karras2019style ; karras2020analyzing techniques, enabling seemingly realistic image generation, but their performance is still unsatisfactory, which limits practical applications. Recently, the diffusion structure blattmann2023stable ; guo2023animatediff ; liu2024towards has greatly enhanced the generation capability of images and videos. Benefiting from this new wave of research, deepfake technology demonstrates potential value for practical applications and can generate content indistinguishable from real ones, which has further attracted attention and is widely applied in numerous fields barattin2023attribute , including entertainment, movie production, online live broadcasting, and privacy protection, etc.
Figure 1: Time diagram that reflects the survey pipeline. Zoom in for a better holistic perception of this work.
@@
Deepfake generation can generally be divided into four mainstream research fields: 1) Face swapping xu2022region ; ancilotto2023ximswap ; shiohara2023blendface is dedicated to executing identity exchanges between two person images; 2) Face reenactment koujan2020head2head ; hsu2022dual ; bounareli2023hyperreenact emphasizes transferring source movements and poses; 3) Talking face generation mir2023dit ; ma2023dreamtalk ; zhang2024dr2 focuses on achieving natural matching of mouth movements to textual content in character generation, and 4) Facial attribute editing sun2022anyface ; huang2023ia ; pang2023dpe aims to modify specific facial attributes of the target image. The development of related foundational technologies has gradually shifted from single forward GAN models goodfellow2014generative ; karras2019style to multi-step diffusion models ho2020denoising ; rombach2022high ; blattmann2023stable with higher quality generation capabilities, and the generated content has also gradually transitioned from single-frame images to temporal video modeling guo2023animatediff . In addition, NeRF mildenhall2020nerf ; gao2022nerf has been frequently incorporated into modeling to improve multi-view consistency capabilities jiang2022nerffaceediting ; zhang2022fdnerf .
@@
While enjoying the novelty and convenience of this technology, the unethical use of it raises concerns over the spread of privacy invasion, and dissemination of fake news, necessitating the development of effective forgery detection methods tan2023learning in opposition, a.k.a., deepfake detection. From the earliest handcrafted feature-based methods he2019detection ; zhou2017two to deep learning-based methods yin2023dynamic ; yang2023masked , and the recent hybrid detection techniques ilyas2023avfakenet , forgery detection has undergone substantial technological advancements along with the development of generative technologies. The data modality has also transitioned from the spatial and frequent domains qian2020thinking ; li2021frequency to the more challenging temporal domain haliassos2022leveraging ; yang2023avoid . Considering that current generative technologies have a higher level of interest, develop faster, and can generate indistinguishable content from reality mirsky2021creation , corresponding detection technologies need continuous evolution.
Overall, despite the significant progress made in both directions, they still exhibit challenging limitations in specific scenarios vyas2024analysing , mainly reflected in the visual perception authenticity and generative accuracy of models. This has attracted a large number of researchers to continue their efforts and has sparked thoughts on industrial applications. Existing survey works only focus on partial deepfake fields and lack discussions on new technologies sha2023deep ; mirsky2021creation ; liu2023gan , especially diffusion-based image/video generation methods, due to their disconnection from current technologies. The survey will comprehensively discuss these fields as well as related sub-fields, and cover tracking of the latest works.
Figure 2: Top: Illustration of different deepfake generation (Sec. 3.1) and detection tasks (Sec. 3.2) that are discussed in this survey. Bottom: Specific facial attribute modification of each task. Data from NVIDIA Keynote at COMPUTEX 2023 at 29:40.
@@
$\bullet$ Contribution. In this survey, we comprehensively explore the key technologies and latest advancements in Deepfakes generation and forgery detection. We first unify the task definitions (Sec. 2.1), provide a comprehensive comparison of datasets and metrics (Sec. 2.3), and discuss the development of related technologies. Specifically, we investigate four mainstream deepfake fields: face swapping (Sec. 3.1.1), face reenactment (Sec. 3.1.2), talking face generation (Sec. 3.1.3), and facial attribute editing (mainly on multiple editing) (Sec. 3.1.4), as well as forgery detection (Sec. 3.2). We also analyze the benchmarks and settings for each domain, thoroughly evaluating the latest and influential works published in top-tier conferences/journals (Sec. 4), especially recent diffusion-based approaches. Additionally, we discuss closely related fields, including head swapping, face super-resolution, face reconstruction, face inpainting, body animation, portrait style transfer, makeup transfer, and adversarial sample detection. Benefiting from the current popularity of AIGC, the research iteration cycle in the deepfake field has been significantly reduced, and we keep updating and discussing the in-submission works in the revised version.
@@
$\bullet$ Scope. This survey primarily focuses on mainstream face-related tasks, including face swapping, face reenactment, talking face generation, facial (multiple) attribute editing, and forgery detection. We also cover some related domain tasks in Sec. 2.4 and detail specific popular sub-tasks in Sec. 3.3. Considering the large number of articles (including published and preprints), we mainly include representative and attention-grabbing works. In addition, we compare this investigation with recent surveys. Sha et alsha2023deep only discuss character generation while we cover a more comprehensive range of tasks. Compared to works mirsky2021creation ; liu2023gan ; melnik2024face , our study encompasses a broader range of technical models, particularly the more powerful diffusion-based methods. Additionally, we thoroughly discuss the related sub-fields of deepfake generation and detection.
$\bullet$ Survey Pipeline. Fig. 1 shows the pipeline of this survey. Sec. 2 provides the essential background knowledge that encompasses task-specific definitions, datasets, evaluation metrics, and exploration of related research areas. Sec. 3 presents technical discussions, examining the four most popular deepfake tasks from the perspective of technological categorization and evolution. In addition, we meticulously classify and discuss forgery detection techniques, emphasizing the technological routes involved. Subsequently, Sec. 4 organizes and evaluates performance from various approaches for a thorough and fair performance comparison. Sec. 5 focuses on a critical review of the challenges that persist in existing techniques, outlining feasible future development directions. Finally, we encapsulate the entirety of the paper through a comprehensive summary in Sec. 6.
2 Background
@@
In this section, we first introduce the conceptual definitions of the discussed mainstream fields. Fig. 2 illustrates the intuitive objectives for each task and shows the distinctions among tasks in terms of manipulated facial components. Then, we review the developmental history of commonly used neural networks, highlighting several representative ones. Next, we summarize popular datasets, metrics, and loss functions. Finally, we comprehensively discuss several relevant domains.
2.1 Problem Definition
@@
where $\phi_{D}$ abstracts the specific detection network and $S_{o}$ represents the fake score for the generated content $I_{o}$.
@@
$\bullet$ Face Swapping involves replacing the identity information of the target face $I_{t}$ by the source face $C$=$I_{s}$, while preserving ID-irrelevant attributes of the target face, such as skin color and expressions.
$\bullet$ Face Reenactment alters the facial movement of a target image $I_{t}$ without changing the identity and other facial attributes, under the conditions of driving image, video, head pose, etc.. This technology often involves the support of facial motion capture technology, such as facial tracking or prediction based on deep learning models zheng2021efficient .
Figure 3: Development timeline of three mainstream generative modals, i.e., VAE, GAN, and Diffusion.
$\bullet$ Talking Face Generation can be viewed as an extension in time, aiming at generating a talking video $I_{o}$=$\{I_{o}^{i}\},i=0,1,\cdots,N-1$ with the character in the target image $I_{t}$ engaging in a conversation based on an arbitrary driving source, such as text, audio, video, or a multi-modal composite source. The lip movements, facial poses, expressions, emotions, and spoken content information of the character in the generated video match the target information.
$\bullet$ Facial Attribute Editing aims to modify the semantic information of the target face $I_{t}$ (e.g., personality, age, expressions, skin color, etc.) in a directed manner based on individual interest and preference. Existing methods include single and comprehensive attribute editing: the former focuses on training a model for only one attribute, while the latter integrates multiple attribute editing tasks simultaneously that is our primary focus in this survey.
@@
$\bullet$ Forgery Detection aims to detect and identify anomalies, tampering, or forgery areas in images or videos by the anomaly score $S_{o}$, and it has great research and application value in information security and multimedia forensics.
2.2 Technological History and Roadmap
$\bullet$ Generative Framework. Variational Autoencoders (VAEs) kingma2014auto ; sohn2015learning ; van2017neural , Generative Adversarial Networks (GANs) goodfellow2014generative ; karras2019style ; karras2020analyzing , and Diffusion sohl2015deep ; ho2020denoising ; rombach2022high have played pivotal roles in the developmental history of generative models.
1) VAE kingma2014auto emerges in 2013, altering the relationship between latent features and linear mappings in autoencoder latent spaces. It introduces feature distributions like the Gaussian distribution and achieves the generation of new entities through interpolation. To enhance generation capabilities under specific conditions, CVAE sohn2015learning introduces conditional input. VQ-VAE van2017neural introduces the concept of vector quantization to improve the learning of latent representations. Subsequent models continually advance.
2) GANs goodfellow2014generative achieve high-quality generation through adversarial training with an extra discriminator. Subsequently, research on GANs experiences a surge, and currently, implementations based on GANs remain the mainstream approach for various deepfake tasks. CGAN mirza2014conditional introduces conditional control variables to GANs. Pix2Pix isola2017image enhances GAN performance in specific image translation tasks. StyleGAN karras2019style introduces the concept of style transfer to GANs. StyleGAN2 karras2020analyzing further improves the quality and controllability of generated images. Additionally, some composite research combines GANs with VAEs, e.g., CVAE-GAN bao2017cvae .
3) Diffusion models sohl2015deep model the generation of data as a diffusion process. DDPM ho2020denoising gains widespread attention for its outstanding generative performance, especially when the model excels in handling large-scale, high-resolution images. LDM rombach2022high is more flexible and powerful when it comes to modeling complex data distributions. In the field of video generation, diffusion models play a crucial role ma2023follow ; liu2023residual ; gupta2023photorealistic . SVD blattmann2023stable fine-tunes the base model using an image-to-video conversion task based on text-to-video models. AnimateDiff guo2023animatediff attaches a newly initialized motion modeling module to a frozen text-to-image model and trains it on subsequent video clips to refine reasonable motion prior knowledge, thus achieving excellent generation results.
@@ 2.2 Technological History and Roadmap
Figure 4: Works summarization on different directions per year. Data is obtained on 2024/05/13.

### v5

@@ Abstract.
Deepfake technology aims to synthesize highly realistic facial images and videos, with broad application potential in entertainment, film production, and digital human modeling. Deep learning has driven major progress in generative modeling, from VAEs and GANs to the recent rise of diffusion models. The latter have sparked a renewed wave of research through their superior generation quality. In addition to deepfake generation, corresponding detection technologies continuously evolve to regulate the potential misuse of deepfakes, such as privacy invasion and phishing attacks. This survey comprehensively reviews the latest developments in deepfake generation and detection, summarizing and analyzing current state-of-the-arts in this rapidly evolving field. First, we unify task definitions, comprehensively introduce datasets and metrics, and summarize the underlying technologies. Then, we review the development of several related sub-fields and examine four representative deepfake research fields: face swapping, face reenactment, talking-face generation, and facial attribute editing, as well as forgery detection. Subsequently, we benchmark representative methods on widely adopted datasets to provide a comprehensive and up-to-date evaluation of the most influential published works. Finally, we discuss the key challenges and outline future research directions for the field. We closely follow the latest developments in this project.
@@
Deepfake Generation, Face Swapping, Face Reenactment, Talking Face Generation, Facial Attribute Editing, Forgery Detection, Survey.
1. Introduction
@@
Artificial Intelligence Generated Content (AIGC) garners considerable attention (sha2023deep, ) in academia and industry. Deepfake generation, as one of the important technologies in the generative domain, gains significant attention due to its ability to create highly realistic facial media content. This technique transitions from traditional graphics-based methods to deep learning-based approaches. Early methods employ advanced Variational Autoencoder (kingma2014auto, ; sohn2015learning, ; van2017neural, ) (VAE) and Generative Adversarial Network (GAN) (karras2019style, ; karras2020analyzing, ) techniques, enabling seemingly realistic image generation, but their performance is still unsatisfactory, which limits practical applications. Recently, the diffusion structure (blattmann2023stable, ; guo2023animatediff, ; liu2024towards, ) has greatly enhanced the generation capability of images and videos. Benefiting from this new wave of research, deepfake technology demonstrates potential value for practical applications and can generate content indistinguishable from real ones, which has further attracted attention and is widely applied in numerous fields (barattin2023attribute, ), including entertainment, movie production, etc.
@@
Deepfake generation can generally be divided into four mainstream research fields: 1) Face swapping (xu2022region, ; ancilotto2023ximswap, ; shiohara2023blendface, ) is dedicated to executing identity exchanges between two person images; 2) Face reenactment (hsu2022dual, ; bounareli2023hyperreenact, ) emphasizes transferring source movements and poses; 3) Talking face generation (mir2023dit, ; zhang2024dr2, ) focuses on achieving natural matching of mouth movements to textual content in character generation, and 4) Facial attribute editing (sun2022anyface, ; huang2023ia, ; pang2023dpe, ) aims to modify specific facial attributes of the target image. The development of related foundational technologies has gradually shifted from single forward GAN models (goodfellow2014generative, ; karras2019style, ) to multi-step diffusion models (ho2020denoising, ; rombach2022high, ; blattmann2023stable, ) with higher quality generation capabilities, and the generated content has also gradually transitioned from single-frame images to temporal video modeling (guo2023animatediff, ). In addition, NeRF (mildenhall2020nerf, ; gao2022nerf, ) has been frequently incorporated into modeling to improve multi-view consistency capabilities (jiang2022nerffaceediting, ; zhang2022fdnerf, ).
@@
While enjoying the novelty and convenience of this technology, its unethical use has raised serious societal and ethical concerns. In practice, deepfake content has been used to generate non-consensual explicit videos targeting individuals. Moreover, there have been instances where videos of deceased public figures were synthetically recreated and disseminated without the consent of their families, causing emotional distress and raising serious moral and legal questions. These incidents illustrate substantial risks of privacy invasion, identity impersonation, and large-scale dissemination of misleading or harmful content. Consequently, there is an urgent need for effective forgery detection systems to mitigate malicious misuse and safeguard information security (li2021image, ; li2024unionformer, ). From the earliest handcrafted feature-based methods (he2019detection, ; zhou2017two, ) to deep learning-based methods (yin2023dynamic, ; yang2023masked, ), and the recent hybrid detection techniques (ilyas2023avfakenet, ), forgery detection has undergone substantial technological advancements along with the development of generative technologies. The data modality has also transitioned from the spatial and frequent domains (qian2020thinking, ; li2021frequency, ) to the more challenging temporal domain (haliassos2022leveraging, ; yang2023avoid, ). Considering that current generative technologies have a higher level of interest, develop faster, and can generate indistinguishable content from reality (mirsky2021creation, ), corresponding detection technologies need continuous evolution.
Overall, despite notable progress in both directions, existing methods still face limitations in visual authenticity and generative accuracy under challenging scenarios (vyas2024analysing, ). These issues continue to draw research attention and raise questions about their practical deployment. However, prior surveys cover only parts of the deepfake landscape and overlook emerging technologies (sha2023deep, ; mirsky2021creation, ; liu2023gan, ), particularly diffusion-based image and video generation. This survey provides a comprehensive overview of these areas and related sub-fields while also tracking the latest developments.
@@
$\bullet$ Contribution. In this survey, we comprehensively explore the key technologies and latest advancements in Deepfakes generation and forgery detection. We first unify the task definitions (Sec. 2.1), provide a comprehensive comparison of datasets and metrics (Sec. 2.3), and discuss the development of related technologies. Then, we investigate four mainstream deepfake fields, as well as forgery detection (Sec. 3). We also analyze the benchmarks and settings for each domain, thoroughly evaluating the latest and influential works published in top-tier conferences/journals (Sec. 4), especially recent diffusion-based approaches. Additionally, we discuss closely related fields, including head swapping, face super-resolution, face reconstruction, face inpainting, body animation, portrait style transfer, makeup transfer, and adversarial sample detection.
@@
$\bullet$ Scope. This survey primarily focuses on four mainstream face-related tasks and forgery detection. We also cover some related domain tasks in Sec. 2.4 and detail specific popular sub-tasks in Sec. 3.3. Considering the large number of articles (including published and preprints), we mainly include representative and attention-grabbing works. In addition, we compare this investigation with recent surveys. Sha et al(sha2023deep, ) only discuss character generation while we cover a more comprehensive range of tasks. Compared to works (mirsky2021creation, ; liu2023gan, ; melnik2024face, ), our study encompasses a broader range of technical models, particularly the more powerful diffusion-based methods. Additionally, we thoroughly discuss the related sub-fields of deepfake generation and detection.
$\bullet$ Survey Methods. To conduct an effective literature review, we systematically searched major academic databases, including IEEE Xplore, ACM Digital Library, Springer, and ScienceDirect, and screened technical articles that met our inclusion criteria. Specifically, we focused on studies published between 2020 and 2025, while grouping research published prior to 2020 into a separate category. For each research topic, we combined relevant technical keywords and application-oriented keywords to perform targeted searches. The technical keywords included Generative Adversarial Networks (GANs), GAN variants, diffusion models, autoencoders, and Neural Radiance Fields (NeRF), among others. The application keywords included deepfake generation, deepfake detection, face swapping, face replacement, face reenactment, talking face, talking head, and forgery detection. The retrieved results were manually deduplicated and categorized according to application domains. Representative works for each research topic were selected based on publication venue quality, citation count, and technical novelty.
Figure 1. Time diagram that reflects the survey pipeline. Zoom in for a better holistic perception of this work.
$\bullet$ Survey Pipeline. Fig. 1 illustrates the overall structure of this survey. Sec. 2 introduces the essential background, including task definitions, datasets, evaluation metrics, and related research areas. It also outlines key applications of deepfake technologies and discusses their ethical implications. Sec. 3 presents a technical review of the four major deepfake tasks, organized from the perspective of methodological categorization and technological evolution. Sec. 4 summarizes and compares the performance of representative methods on widely used benchmarks to provide a fair and comprehensive evaluation. Sec. 5 critically examines remaining challenges and highlights potential directions for future research. Finally, Sec. 6 provides a concise summary of the survey.
2. Background
@@
In this section, we first introduce the conceptual definitions of the discussed mainstream fields. Fig. 2 illustrates the intuitive objectives for each task and shows the distinctions among tasks in terms of manipulated facial components. Then, we review the developmental history of commonly used neural networks, highlighting several representative ones. Next, we summarize popular datasets, metrics, and loss functions. Subsequently, we comprehensively discuss several relevant domains. Finally, we introduce some application and discussion about there ethical considerations.
2.1. Problem Definition
@@
where $\phi_{D}$ abstracts the detection network and $S_{o}$ represents the fake score for generated content $I_{o}$.
@@
$\bullet$ Face Swapping. This task replaces the identity of the target face $I_{t}$ with that of the source face $I_{s}$, while maintaining target-specific, ID-irrelevant attributes such as skin tone and expressions.
$\bullet$ Face Reenactment. This task transfers the facial movements from a driving image or video to a target image $I_{t}$, while keeping the target’s identity and attributes unchanged. It commonly relies on facial motion capture techniques, including tracking or deep-learning-based motion prediction.
$\bullet$ Talking Face Generation. This task generates a talking video $I_{o}$ for the character in a target image $I_{t}$, driven by text, audio, video, or multimodal inputs. The output should accurately reflect the driving information, including lip motion, facial pose, emotions, and spoken content.
Figure 2. Top: Illustration of different deepfake generation and detection tasks that are discussed in this survey. Bottom: Specific facial attribute modification of each task. Data from NVIDIA Keynote at COMPUTEX 2023.
$\bullet$ Facial Attribute Editing. This task modifies semantic attributes of a target face $I_{t}$ (e.g., age, expression, or skin tone) in a controlled way according to user intent. Methods fall into single-attribute and multi-attribute editing, with the latter is the main focus of this survey.
@@
$\bullet$ Forgery Detection. This task detects and localizes tampering or forged regions in images or videos using an anomaly score $S_{o}$. It is importance to information security and multimedia forensics.
Figure 3. Development timeline of three mainstream generative modals, i.e., VAE, GAN, and Diffusion.
@@
Figure 4. Works summarization on different directions per year. Data is obtained on 2025/11/20.
2.2. Technological History and Roadmap
$\bullet$ Generative Framework. VAEs (kingma2014auto, ; sohn2015learning, ; van2017neural, ), GANs (goodfellow2014generative, ; karras2019style, ; karras2020analyzing, ), and Diffusion (sohl2015deep, ; ho2020denoising, ; rombach2022high, ) have played pivotal roles in the developmental history of generative models.
1) VAE (kingma2014auto, ) emerges in 2013, altering the relationship between latent features and linear mappings in autoencoder latent spaces. It introduces feature distributions like the Gaussian distribution and achieves the generation of new entities through interpolation. To enhance generation capabilities under specific conditions, CVAE (sohn2015learning, ) introduces conditional input. VQ-VAE (van2017neural, ) introduces the concept of vector quantization to improve the learning of latent representations.
2) GANs (goodfellow2014generative, ) generate realistic data through an adversarial process between two networks: a generator and a discriminator. This relationship can be likened to a boxing match in which the generator acts as the challenger attempting to produce increasingly convincing fake samples, while the discriminator serves as the defending champion distinguishing real from generated content. As the two opponents improve through continual competition, the quality of generated outputs steadily increases. Built upon this adversarial paradigm, GAN-based methods have rapidly expanded and remain central to many deepfake tasks. Extensions can be viewed as introducing new strategies to this “match.” CGAN (mirza2014conditional, ) conditions both networks on auxiliary information, making the competition more guided. Pix2Pix (isola2017image, ) specializes the framework for image-to-image translation, akin to training the challenger for specific techniques. StyleGAN (karras2019style, ) introduces style-based controls for fine-grained feature manipulation, and StyleGAN2 (karras2020analyzing, ) further stabilizes training to enhance quality and controllability. Hybrid models such as CVAE-GAN (bao2017cvae, ) enrich the generator’s latent modeling by integrating VAE-based representations with adversarial learning.
3) Diffusion models (sohl2015deep, ) treat data generation as a gradual denoising process. Starting from near-random noise, the model learns to reverse the corruption step by step, eventually recovering a clean and realistic image. DDPM (ho2020denoising, ) popularizes this paradigm by achieving outstanding generative performance, particularly on large-scale and high-resolution images. LDM (rombach2022high, ) improves efficiency and flexibility by performing the denoising process in a compact latent space, similar to restoring the compressed negative of an image rather than the full-resolution photograph. SVD (blattmann2023stable, ) fine-tunes a base text-to-video model for image-to-video conversion, akin to extending a single frame into a coherent sequence. AnimateDiff (guo2023animatediff, ) attaches a motion modeling module to a frozen text-to-image backbone and trains it on short video clips.
$\bullet$ Discriminative Neural Network. Convolutional Neural Networks (CNNs) (lecun1998gradient, ; krizhevsky2012imagenet, ; he2016deep, ; liu2022convnet, ; emo, ; eatformer, ) have played a pivotal role in the history of deep learning. LeNet (lecun1998gradient, ), as the pioneer of CNNs, showcased the charm of machine learning. AlexNet (krizhevsky2012imagenet, ) and ResNet (he2016deep, ) made deep learning feasible. Recently, ConvNeXt (liu2022convnet, ) has achieved excellent results surpassing those of Swin-Transformer (liu2021swin, ). The Transformer architecture initially proposed (vaswani2017attention, ) in 2017. The core idea involves using self-attention mechanisms to capture dependencies between different positions in the input sequence, enabling global modeling of sequences. ViT (dosovitskiy2020image, ) demonstrates that using Transformer in the field of computer vision can still achieve excellent performance. In addition, Swin-Transformer (liu2021swin, ) addresses the limitations of Transformer in handling high-resolution image processing tasks. Swin-Transformer V2 (liu2022swin, ) further improves the model’s efficiency and the resolution of manageable inputs.
$\bullet$ Neural Radiance Field. NeRF, introduced in 2020 (mildenhall2020nerf, ), leverages volume rendering and implicit neural fields to represent and reconstruct the geometry and illumination of 3D scenes (gao2022nerf, ). Compared to traditional 3D methods, it exhibits higher visual quality and is currently widely applied in tasks such as 3D geometry enhancement (sharma2023neural, ; irshad2023neo, ), segmentation (liu2023instance, ) and 6D pose estimation (min2023neurocs, ). In addition, Some notable works (hwang2023faceclipnerf, ; yang2023urbangiraffe, ) combining NeRF as a supplement to 3D information and generation models are particularly prominent at present.
2.3. Datasets, Metrics, and Losses
$\bullet$ Dataset. Given the various datasets in surveyed fields, we use numerical labels to save post-textual space. 1) Commonly used deepfake generation datasets include LFW (huang2008labeled, ), CelebA (liu2015deep, ), CelebA-HQ (karras2018progressive, ), VGGFace (parkhi2015deep, ), VGGFace2 (cao2018vggface2, ), FFHQ (karras2019style, ), Multi-PIE (moore2010multi, ), VoxCeleb1 (nagrani2017voxceleb, ), VoxCeleb2 (chung2018voxceleb2, ), MEAD (wang2020mead, ), MM CelebA-HQ (xia2021tedigan, ), CelebAText-HQ (sun2021multi, ), CelebV-HQ (zhu2022celebv, ), TalkingHead-1KH (wang2021one, ), LRS2 (afouras2018deep, ), LRS3 (afouras2018lrs3, ), etc. 2) Commonly used forgery detection datasets include UADFV (li2018ictu, ), DeepfakeTIMIT (korshunov2018deepfakes, ), FF++ (rossler2019faceforensics++, ), Deeperforensics-1.0 (jiang2020deeperforensics, ), DFDCP (dolhansky2019deepfake, ), DFDC (dolhansky2020deepfake, ), Celeb-DF (li2020celeb, ), Celeb-DFv2 (li2020celeb, ), FakeAVCeleb (khalid2021fakeavceleb, ), DFD (DfD, ), WildDeepfake (zi2020wilddeepfake, ), KoDF (kwon2021kodf, ), UADFV (yang2019exposing, ), Deephy (narayan2022deephy, ), DF-Platter (narayan2023df, ), etc. We summarize popular datasets in Tab. 1 .
$\bullet$ Metric. 1) For deepfake generation tasks, commonly used metrics include: Peak Signal-to-Noise Ratio (PSNR) (wang2023memory, ), Structured Similarity (SSIM) (wang2004image, ), Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable, ), Fréchet Inception Distance (FID) (heusel2017gans, ), Kernel Inception Distance (KID) (binkowski2018demystifying, ), Cosine Similarity (CSIM) (wang2021one, ), Identity Retrieval Rate (ID Ret) (wang2018cosface, ), Expression Error (chaudhuri2019joint, ), Pose Error (ruiz2018fine, ), Landmark Distance (LMD) around the mouths (chen2018lip, ), Lip-sync Confidence (LSE-C) (prajwal2020lip, ), Lip-sync Distance (LSE-D) (prajwal2020lip, ), etc. 2) For forgery generation commonly uses: Area Under the ROC Curve (AUC) (le2023quality, ), Accuracy (ACC) (wang2022m2tr, ), Equal Error Rate (EER) (huang2023implicit, ), Average Precision (AP) (tan2023learning, ), F1-Score (chen2021image, ), etc. Detailed definitions are explained in Sec. 4.1.
$\bullet$ Loss Function. VAE-based approaches generally employ reconstruction loss and KL divergence loss (kingma2014auto, ). Commonly used reconstruction loss functions include Mean Squared Error, Cross-Entropy, LPIPS (zhang2018unreasonable, ), and perceptual (johnson2016perceptual, ) losses. GAN-based methods further introduce adversarial loss (goodfellow2014generative, ) to increase image authenticity, while diffusion-based works introduce denoising loss function (ho2020denoising, ).